OpenStudy (anonymous):

A train is moving at a velocity of 120 km/h. Find the distance will it travel in 30 s.

OpenStudy (anonymous):

@AravindG plzz help

OpenStudy (anonymous):

@mathslover

OpenStudy (aravindg):

velocity =distance/time :)

OpenStudy (anonymous):

okay then

mathslover (mathslover):

first of all : \[\huge{\mathbb{WElcome}\textbf{To}\mathbb{OPen}\textbf{Study}}\]

OpenStudy (aravindg):

velocity =siplacement /time is better

OpenStudy (anonymous):

okay... thanks

OpenStudy (anonymous):

kk

OpenStudy (aravindg):

displacement/time

mathslover (mathslover):

given velocity = 120 km/h = 120 * 5/18 m/s = 600/18 m/s

OpenStudy (anonymous):

can u give me explained solution

OpenStudy (anonymous):

33.33

mathslover (mathslover):

\[\huge{velocity=\frac{displacement}{time}}\] \[\huge{\frac{600}{18}=\frac{displacement}{30}}\] \[\huge{\frac{600}{18}*30 = displacement}\] \[\huge{1000\textbf{meter}=displacement}\]

OpenStudy (anonymous):

can t get u

mathslover (mathslover):

where r u having problem ?

OpenStudy (anonymous):

a detailed explaination on it + solution

mathslover (mathslover):

see do u know velocity = displacement / time ?

OpenStudy (anonymous):

k

mathslover (mathslover):

now we have velocity = 120 km/h right ?

OpenStudy (anonymous):

correct i get u now

mathslover (mathslover):

r u sure that u got it ?

OpenStudy (anonymous):

yes

mathslover (mathslover):

that is great !!! best of luck for your journey in openstudy .... just a brief introduction to u of this site : \( \begin{array}{} \cal{\large{\text{Hey! Good to see you on this site!}}}\\ \\ {\huge \mathtt{\color{blue}{\text{Welcome To OpenStudy!}}}} \\ \text{enjoy answering and asking questions here}\\ \cal{\large{\text{Tips:)}}} \\ \textbf{Ambassadors (people with A next to their name) are here to help you.}\\ \textbf{If you have any questions about the site, feel free to ask us!}\\ \textbf{Users who appear in Purple are Moderators and help run the site.}\\ \textbf{If you have a serious issue with the site or with the behavior of other}\\ \textbf{users}\\ \textbf{please feel free to raise it with them.}\\ \textbf{Want to study a particular subject? Click 'Find More Subjects'}\\ \textbf{search for it there.}\\ \textbf{If you get a good response from a user , click on best response}\\ \end{array}\) go to this link for more information : http://www.openstudy.com/code-of-conduct